OpenStudy (anonymous):

i have to calculate the angle between v and w. v= (1, 3, 7) w= -3x+2y+2z so w= (-3, 2, 2). so how do i find the angle between these two guys?

OpenStudy (anonymous):

by the way, these are vectors we're working with here.

OpenStudy (zarkon):

\[v\cdot w=||{v}||\,||w||\cos(\theta)\]

OpenStudy (anonymous):

their lengths are IIvII=root 59 and IIwII=root 17, but i get lost after that part

OpenStudy (zarkon):

can you get \[v\cdot w\]?

OpenStudy (anonymous):

argh. if i multiply them together its root 1003, but that makes me feel like i did something wrong. maybe the lengths i calculated were wrong..

OpenStudy (zarkon):

dot product...you should get 17

OpenStudy (zarkon):

\[1*(-3)+3*2+7*2\]

OpenStudy (anonymous):

i did cross product, idk why. okay so when you get 17, what do you do next?

OpenStudy (zarkon):

plug in all the info into the equation I gave above...solve for \[\theta\]

OpenStudy (zarkon):

\[\theta=\cos^{-1}\left(\frac{v\cdot w}{\|v\| \|w\|}\right)\]

OpenStudy (anonymous):

θ=cos−1(17/root59 root17) woo hoo! is that good?

OpenStudy (zarkon):

yes...which is about 1.004 radians or 57.535 deg

OpenStudy (anonymous):

you're the best zarkon!

OpenStudy (zarkon):

don't feed my ego ;)